OpenStudy (anonymous):

How do you solve... Susan Jacobson has grades of 76 and 81 on her first to algebra tests. If she wants an average of at least 80 after her third test, what score must she make on her test?

OpenStudy (ash2326):

@melb94 Do you know how to find average of three no.s?

OpenStudy (anonymous):

no i do not

OpenStudy (ash2326):

Do you understand what average implies?

OpenStudy (ash2326):

@melb94

OpenStudy (anonymous):

average would be the sum of all the test?

OpenStudy (ash2326):

Sum of all the test marks is the Sum, Suppose you take a series of exams and you score different in each of the exam, suppose we want to know how well you performed, we would know that from the average. Suppose you scored 90, 95, 100, 80, 85 Do you follow this?

OpenStudy (anonymous):

I think so.

OpenStudy (ash2326):

To find average we'll find the sum of all the test scores S=90+95+100+80+85 Can you find the sum?

OpenStudy (anonymous):

450

OpenStudy (ash2326):

Good and how many exam did you take?

OpenStudy (anonymous):

5

OpenStudy (ash2326):

Good then you average score would be \[A=\frac{S}{5}=\frac{450}{5}=\] Can you find this?

OpenStudy (anonymous):

90?

OpenStudy (ash2326):

yes, that's how average is calculated. Did you follow?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

ok, let's work on your problem. Can you try?

OpenStudy (anonymous):

So the equation would be?\[76+81+x \ge80\]

OpenStudy (ash2326):

yes, that's how average is calculated. Did you follow?

OpenStudy (ash2326):

We will use equal to symbol and you have to make use of the no. of tests

OpenStudy (anonymous):

so 76+81+80/3?

OpenStudy (ash2326):

but we don't know the score of the third test, let it be x

OpenStudy (anonymous):

so 76+81+x=157+x and then divide by 3?

OpenStudy (ash2326):

yes, and we already have the average in the question

OpenStudy (anonymous):

so the answer would be 83?

OpenStudy (ash2326):

yes, good work.

OpenStudy (anonymous):

yay! thanks

OpenStudy (ash2326):

Welcome :D